A 200-gram solution is 25% acid. How much pure acid must be added to produce a solution that is 40% acid?
Answer by checkley77(12844)

.25*200+X=.40(200+X)
50+X=80+.40X
X-.4X=80-50
.6X=30
X=30/.6
X=50 GRAMS OF PURE ACID IS USED.
PROOF:
.25*200+50=.40(200+50)
50+50=.40*250
100=100